OpenStudy (anonymous):

solve the following equations on (0 less than or equal to theta less than or equal to 2 pi) 1.)tan 2 theta=1 2.)sin 2theta=square root 2costheta for #1 i get pi/3 and 5pi/3 and for #2 pi/4 but am not sure if theye are correct

OpenStudy (zzr0ck3r):

1) let x = cos(theta) 2x^2+x=0 x(2x+1) = 0 so cos(theta) = 0 and 2cos(theta) + 1 = 0 ---> cos(theta) = -1/2 so solve these cos(theta) = 0 and cos(theta) = -1/2 between 0 and 2pi this is true when theta = pi/2, 3pi/2, 4pi/6, and 8pi/6

OpenStudy (campbell_st):

so is the 1st question 1. tan(2x) = 1 or \[2. ... \tan^2(x) = 1\]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

i made a mistake when i wrote the answers i got for that problem with another problem that had a tan theta in it but the pi/8 is what i got

OpenStudy (campbell_st):

ok... so its easy so tan is positive in the 1st and 3rd quadrants and you know \[2x = \tan^{-1}(1) .......or..... 2x = \frac{\pi}{4}\] so find x 1st quadrant the the answer you find. 3rd quadrant is \[\pi + answer\]

OpenStudy (campbell_st):

hope that helps

OpenStudy (campbell_st):

well you are correct with 1 answer, but there are 2 in the given domain.,

OpenStudy (campbell_st):

and for the 2nd equation is it \[\sin(2x) = \sqrt{2}\cos(x)\]

OpenStudy (anonymous):

yeah thats the second equation

OpenStudy (anonymous):

so u say 2x=pi/4 wud i divide both sides by 2 to get just x giving me 1/8?

OpenStudy (anonymous):

and then i just add pi to 1/8?

OpenStudy (campbell_st):

well actually \[x = \frac{\pi}{8}\] the other answer is \[x = \pi + \frac{\pi}{8}....or......x = \frac{\pi}{8}\]

OpenStudy (campbell_st):

opps should be \[x = \frac{9\pi}{8}\]

OpenStudy (campbell_st):

so 2 answers to the question

OpenStudy (anonymous):

ohh my bad yeah i ment pi/8 but i just didnt put the pi sorry

OpenStudy (campbell_st):

so looking at the 2nd problem sin(2x) = 2sin(x)cos(x) so you need this substitution for the 2nd equation..

OpenStudy (campbell_st):

so its \[2 \sin(x)\cos(x) = \sqrt{2}\cos(x)\] divide both sides by 2cos(x) gives \[\sin(x) = \frac{\sqrt{2}}{2}\] this is an exact value. so \[x = \frac{\pi}{4}\] again because of the domain. there are 2 solutions 1st quadrant, as shown above and 2nd quadrant which uses \[\pi - answer\] hope this helps

OpenStudy (anonymous):

so 3pi/4 is the second answer?

OpenStudy (campbell_st):

I'd say so

OpenStudy (zzr0ck3r):

so confused...did the question get edited and changed? Or did I just make up a problem and solve it?

OpenStudy (anonymous):

no lol i edited the question n removed the one u had helped me solve

OpenStudy (anonymous):

thank for your help tho =D ... i have final thursday on doing problems like this n im so confused on how to do it n i gotta pass or else i fail the class since i didnt do so well on the test =[

OpenStudy (zzr0ck3r):

gl